OpenStudy (anonymous):

Write the equation of a circle with a center at (-5, -1) and a radius of 5 Answers: (x + 5)2 + (y + 1)2 = 5 (x + 5)2 + (y + 1)2 = 25 (x – 5)2 + (y – 1)2 = 25 (x – 5)2 + (y – 1)2 = 5

jimthompson5910 (jim_thompson5910):

Hint: The general equation of a circle (x-h)^2+(y-k)^2=r^2 has center: (h,k) radius: r

OpenStudy (anonymous):

Help explain please @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so for example, (x – 5)^2 + (y – 1)^2 = 25 is in the form (x-h)^2+(y-k)^2=r^2 where h = 5 k = 1 r = 5 So we can say Center: (5,1) Radius: 5

OpenStudy (anonymous):

Oh ok Thank You!

OpenStudy (anonymous):

I know its the same thing but I was also having trouble with this one as well @jim_thompson5910

OpenStudy (anonymous):

Write the equation of a circle with a center at (1, 2) and a radius of 1.3.

jimthompson5910 (jim_thompson5910):

this time h = 1 k = 2 r = 1.3 plug them into (x-h)^2+(y-k)^2=r^2 to get your answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do you plug in for x and y @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you leave them as they are